there are 180 multiple choice question in test.a candidate gets 4 marks for every correct ans and for every unattempted or wrongly answerd question one mark is deducted from total score of correct answer.it a candidate scired 450 marks in the test.how many question did he sns correctly.
Answer by Edwin McCravy(20077)

let x = the number of questions answered correctly, then:
180-x = the number of unattempted or wrongly answered questions.

Points earned for the x questions answered correctly = 4x
Points deducted from the 180-x unattempted or wrongly
      answered questions = -1(180-x).

4x - 1(180-x) = 450

 4x - 180 + x = 450
   
     5x - 180 = 450

           5x = 630

            x = 126

Number of questions answered correctly = 126
number of unattempted or wrongly answered questions = 180-126 = 54

Checking 

4×126 = 504 for 126 questions answered correctly
-1×54 = -54 for 54 unattempted or wrongly answered questions
------------------------------------------------------------
       450 = his score  (that checks!)

Edwin